OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates ordered pair 3 comma 5pi/3.

OpenStudy (anonymous):

r = 3, theta = 5pi/3 x = rcos(theta) = 3*cos(5pi/3) = 3*cos(-pi/3) = 3/2 y = rsin(theta) = 3*sin(5pi/3) = 3*sin(-5pi/3) = -3sqrt(3)/2 so you get (3/2, -3sqrt(3)/2)

OpenStudy (anonymous):

Thanks(: